OpenStudy (antonio_xx2):

Find the standard form of the equation of the parabola with a focus at (0, -10) and a directrix at y = 10.

OpenStudy (antonio_xx2):

@kirbykirby

OpenStudy (kaylee_td):

Please post this in the relevant section :)

OpenStudy (antonio_xx2):

wait wut?

OpenStudy (kaylee_td):

This is called "openstudy feedback" not "Mathematics"

OpenStudy (antonio_xx2):

@Kaylee_td and i hope u become 1 and change the world;)

OpenStudy (kaylee_td):

Haha Thanks ;)

OpenStudy (antonio_xx2):

and @Kaylee_td we should talk some time:)

OpenStudy (antonio_xx2):

u can stop now @kirbykirby never mind

OpenStudy (kirbykirby):

So you have in general for the parabola: \((x-h)^2 = 4p(y-k)\) The focus is represented by: \((h, k+p)\) And now to determine the sign of \(p\), you notice that y=10 is the directrix, and the focus is at (0, -10), so the focus is below the directrix, meaning you should have a parabola having this shape \(\cap\), so \(p\) is negative Focus: \((h, k+p))=(0, -10)\), so \(h=0\) and \(k+p =-10\) Now the vertex-focus distance is the same as the vertex-directrix distance, so it should be halfway. And so the you get that k =0, so p=-10 \((x-0)^2=4(-10)(y-0)\) \(x^2=-40y\)